OpenStudy (anonymous):

Gabrielle is 15 years younger than Mikhail. The sum of their ages is 75 . What is Mikhails age?

OpenStudy (shamil98):

Let's use variables and make an equation. let gabrielle be g, and mikhail be m. M + G = 75 M = G + 15

OpenStudy (mertsj):

Let x = Mikhail's age x-15 = Gabrielle's age Add them up and get 75

OpenStudy (shamil98):

Using the system, G = M - 15 as well. you want M's age. so.. M + M - 15 = 75 2m = 90 m = 45